Determine the slope of the line passing through each pair of points
A(5,2), B(7,12)
P(-1,4) Q(5,-5)

Answer by Cromlix(4381)

Using equation:
y2 - y1/x2 - x1
12 - 2/7 - 5
= 10/2
= 5
-5 - 4/ 5 - (-1)
-9/6
= -3/2
Hope this helps.
:-)